# Synesthesia of Machines (SoM)-Aided FDD Precoding with Sensing Heterogeneity: A Vertical Federated Learning Approach

### v1

@@ Abstract
FDD, precoding, heterogeneous multi-modal sensing, vertical federated learning.
@@ I Introduction
In this work, we propose a heterogeneous multi-vehicle, multi-modal (H-MVMM) sensing aided precoding method, aiming to enhancing system sum rate with reduced complexity and pilot overhead. We consider a crucial fact that vehicles in practical applications are equipped with different sensors, allowing them to capture unique communication environment features from different viewpoints. To fully leverage the diversity of the multi-view features, we adopt the vertical FL (VFL) framework since VFL is tailored for situations where users have distinct feature spaces but share the same sample space (i.e, BS precoding). A benefit offered by VFL is the reduction in computational complexity. This is achieved by dispersing the centralized precoding optimization from BS to distributed users, who only need to compute their own precoding vectors based on low-dimensional inputs. To bridge the substantial gap between multi-modal sensing and precoding, we customize three unique data preprocessing methods to transform raw sensory data into precoding-related informative representations guided by Synesthesia of Machines (SoM) [16]. To mitigate the effects of data heterogeneity on VFL training, we design a customized loss function that can accelerate local model convergence. Additionally, we devise an online training strategy based on VFL framework, enabling the proposed scheme to dynamically adjust to user number changes.
@@ II System Model and Problem Formulation
$\vspace{-0.5em}\mathbf{x}=\sum^{K}_{k=1}\mathbf{v}_{k}s_{k}=\mathbf{V}\mathbf{s},$ (1)
[…]
@@
To design $\mathbf{V}$, BS needs to acquire instantaneous downlink CSI, which is conventionally achieved by downlink training and feedback. In our scheme, inspired by [11], we propose to bypass explicit channel estimation at the users and directly design the precoder based on downlink received signals since the mutual information between the received pilot and actual channel is larger than that between estimated channel and actual one. In downlink training stage, BS broadcasts pilot sequence $\mathbf{X}\in\mathbb{C}^{N_{\rm t}\times L_{\rm p}}$ of length $L_{\rm P}$ and the received signals $\mathbf{y}_{k}\in\mathbb{C}^{L_{\rm p}}$ at the $k$-th user can be modeled as:
@@
where $\mathbf{z}_{k}\in\mathbb{C}^{L_{\rm p}}\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I})$ is the additive white Gaussian noise and the pilots transmitted at the $l$-th time instant are subject to the transmit power budget $P$ as $\|\mathbf{x}_{l}\|_{2}^{2}\leq P$. After the $k$-th user obtain its precoding vector $\mathbf{v}_{k}$ by its local NN, it subsequently feeds back $\mathbf{v}_{k}$ to BS in form of $B$ bits as: $\mathbf{b}_{k}=\mathcal{F}(\mathbf{v}_{k})$, where function $\mathcal{F}:\mathbb{C}^{N}\rightarrow\{\pm 1\}^{B}$ represents a certain quantization scheme.
@@
s.t. $\mathbf{b}_{k}=\mathcal{F}(\mathbf{v}_{k})=\mathcal{F}(\mathcal{G}_{k}(\mathbf{y}_{k},\mathbb{M}_{k})),~{}~{}\forall k$ (4b)
@@
where $\mathcal{G}_{k}(\cdot)$ denotes the local NN model of the $k$-th vehicle parameterized by $\mathbf{\Theta}_{k},k=1,\cdots,K$. Three widely-used sensors are considered in this work: Global Positioning System (GPS), red-green-blue (RGB) cameras, and Light Detection and Ranging (LiDAR). $\mathbb{M}_{k}$ denotes the set of informative representations obtained from the $k$-th vehicle’s available sensory data, i.e. $\mathbb{M}_{k}=\{\mathbf{x}_{k}^{\rm GPS},\mathbf{x}_{k}^{\rm RGB},\mathbf{x}_{k}^{\rm LiDAR}\}$ if the $k$-th vehicle is equipped with three types of sensors. For simplicity, all subscripts $k$ denote the $k$-th vehicle and will not be restated in the following content. The optimization objective of this work is the parameter sets of all local models. Fig. 1 depicts the schematic of proposed scheme.
Figure 1: An illustration of the system model.
III Proposed SoM-Aided FDD Precoding Under VFL Framework
@@ III-A1 GPS Preprocessing
Intuitively, GPS indicates the rough location of receivers, providing the LoS component information for BS. However, GPS data is typically presented in coordinate form, which does not align with the data types NNs excel at learning, such as encoded feature vectors. Additionally, BS precoding requires relative angle information between the transceiver rather than the absolute user positions. To address these issues, we devise the following preprocessing scheme tailored for GPS data, as shown in Fig. 2.
@@ III-A3 LiDAR Preprocessing
Figure 4: Processing flow of LiDAR point clouds.
@@ III-B Local Training at Vehicles
$\mathbf{v}_{k}=\mathcal{G}^{\rm I}(\mathbf{m}_{k}^{\rm I}),\enspace\enspace\mathcal{G}^{\rm I}:\mathbb{R}^{L_{\rm G}+L_{\rm R}+L_{\rm L}+L_{\rm S}}\mapsto\mathbb{R}^{2N_{\rm t}}.$ (14)
@@ III-C Training Procedure of the Proposed VFL-Based Precoding
2:  for each iteration $j=1,2,\cdots$ do do
@@
4:        Compute $\mathbf{v}_{k}=\mathcal{G}_{k}(\mathbf{y}_{k},\mathbb{M}_{k})$ and feed it back to BS;

### v2

@@ Abstract
FDD, precoding, heterogeneous multi-vehicle multi-modal sensing, vertical federated learning.
@@ I Introduction
In this work, we propose a heterogeneous multi-vehicle, multi-modal (H-MVMM) sensing aided precoding method, aiming to enhance system sum rate with reduced complexity and pilot overhead. We consider a crucial fact that vehicles in practical applications are equipped with different sensors, allowing them to capture unique communication environment features from different viewpoints. To fully leverage the diversity of the multi-view features, we adopt the vertical FL (VFL) framework since VFL is tailored for situations where users have distinct feature spaces but share the same sample space (i.e, BS precoding). A benefit offered by VFL is the reduction in computational complexity. This is achieved by dispersing the centralized precoding optimization from BS to distributed users, who only need to compute their own precoding vectors based on low-dimensional inputs. To bridge the substantial gap between multi-modal sensing and precoding, we customize three unique data preprocessing methods to transform raw sensory data into precoding-related informative representations guided by Synesthesia of Machines (SoM) [16]. To mitigate the effects of data heterogeneity on VFL training, we design a customized loss function that can accelerate local model convergence. Additionally, we devise an online training strategy based on VFL framework, enabling the proposed scheme to dynamically adjust to user number changes.
@@ II System Model and Problem Formulation
$\mathbf{x}=\sum^{K}_{k=1}\mathbf{v}_{k}s_{k}=\mathbf{V}\mathbf{s},$ (1)
where $\mathbf{v}_{k}\in\mathbb{C}^{N}$ represents the precoding vector for the $k$-th user and is also the $k$-th coloumn of the precoding matrix $\mathbf{V}\in\mathbb{C}^{N\times K}$. The precoding matrix $\mathbf{V}$ satisfies the transmit power constraint $P$ as $\operatorname{Tr}(\mathbf{V}\mathbf{V}^{H})\leq P$. The power of data symbols is also normalized so that $\mathbb{E}[\mathbf{s}\mathbf{s}^{H}]=\mathbf{I}$.
@@
To design $\mathbf{V}$, BS needs to acquire instantaneous downlink CSI, which is conventionally achieved by downlink training and feedback. In our scheme, inspired by [11], we propose to bypass explicit channel estimation at the users and directly design the precoder based on downlink received signals since the mutual information between the received pilot and actual channel is larger than that between estimated channel and actual one. In downlink training stage, BS broadcasts pilot sequence $\mathbf{X}\in\mathbb{C}^{N\times L_{\rm p}}$ of length $L_{\rm P}$ and the received signals $\mathbf{y}_{k}\in\mathbb{C}^{L_{\rm p}}$ at the $k$-th user can be modeled as:
@@
where $\mathbf{z}_{k}\in\mathbb{C}^{L_{\rm p}}\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I})$ is the additive white Gaussian noise and the pilots transmitted at the $l$-th time instant are subject to the transmit power budget $P$ as $\|\mathbf{x}_{l}\|_{2}^{2}\leq P$. After the $k$-th user obtain its precoding vector $\mathbf{v}_{k}$ by its local NN, it subsequently feeds back $\mathbf{v}_{k}$ to BS in form of $Q$ bits as: $\mathbf{b}_{k}=\mathcal{F}(\mathbf{v}_{k})$, where function $\mathcal{F}:\mathbb{C}^{N}\rightarrow\{\pm 1\}^{Q}$ represents a certain quantization scheme [^1] The focus of this work is on how to streamline FDD precoding design and ensure high sum rate performance using heterogeneous multi-vehicle multi-modal sensing. Therefore, for the feedback of precoding vector $\mathbf{v}_{k}$, we adopt a $B$-bit uniform quantization strategy where a floating-point number is represented by $B$ bits..
Figure 1: An illustration of the system model.
@@
s.t. $\mathbf{b}_{k}=\mathcal{F}(\mathbf{v}_{k})=\mathcal{F}(\eta(\mathcal{G}_{k}(\mathbf{y}_{k},\mathbb{M}_{k}))),~{}~{}\forall k$ (4b)
@@
where $\mathcal{G}_{k}(\cdot)$ denotes the local NN model of the $k$-th vehicle parameterized by $\mathbf{\Theta}_{k},k=1,\cdots,K$. The operator $\eta:\mathbb{R}^{2N}\mapsto\mathbb{C}^{N}$ defines a mapping that transforms a real-valued vector into a complex-valued vector by taking the first $N$ elements as the real component and the remaining $N$ elements as the imaginary component. Three widely-used sensors are considered in this work: Global Positioning System (GPS), red-green-blue (RGB) cameras, and Light Detection and Ranging (LiDAR). $\mathbb{M}_{k}$ denotes the set of informative representations obtained from the $k$-th vehicle’s available sensory data, i.e. $\mathbb{M}_{k}=\{\mathbf{x}_{k}^{\rm GPS},\mathbf{x}_{k}^{\rm RGB},\mathbf{x}_{k}^{\rm LiDAR}\}$ if the $k$-th vehicle is equipped with three types of sensors. For simplicity, all subscripts $k$ denote the $k$-th vehicle and will not be restated in the following content. The optimization objective of this work is the parameter sets of all local models. Fig. 1 depicts the schematic of proposed scheme.
III Proposed H-MVMM Precoding Scheme Under VFL Framework
@@ III-A1 GPS Preprocessing
Intuitively, GPS indicates the rough location of receivers, providing the line-of-sight (LoS) information for BS. However, GPS data is typically presented in coordinate form, which does not align with the data types NNs excel at learning, such as encoded feature vectors. Additionally, BS precoding requires relative angle information between the transceiver rather than the absolute user positions. To address these issues, we devise the following preprocessing scheme tailored for GPS data, as shown in Fig. 2.
@@ III-A3 LiDAR Preprocessing
Figure 4: Processing flow of LiDAR point clouds.
@@ III-B Local Training at Vehicles
$\mathbf{v}_{k}=\eta(\mathcal{G}^{\rm I}(\mathbf{m}_{k}^{\rm I})),\enspace\enspace\mathcal{G}^{\rm I}:\mathbb{R}^{L_{\rm G}+L_{\rm R}+L_{\rm L}+L_{\rm S}}\mapsto\mathbb{R}^{2N_{\rm t}}.$ (14)
@@ III-C Training Procedure of the Proposed VFL-Based Precoding
2:  for each iteration $j=1,2,\cdots$  do
@@
4:        Compute $\mathbf{v}_{k}=\eta(\mathcal{G}_{k}(\mathbf{y}_{k},\mathbb{M}_{k}))$ and feed it back to BS;